OpenStudy (anonymous):

solve the equation. √x-5 =4

OpenStudy (anonymous):

x = 81

OpenStudy (anonymous):

thanks! @Shay17

OpenStudy (anonymous):

yw

OpenStudy (isaiah.feynman):

Vagueness..

OpenStudy (anonymous):

what do u mean? @Isaiah.Feynman

OpenStudy (isaiah.feynman):

Is the square root covering the x alone?

OpenStudy (anonymous):

no it isnt its covering x-5

jagr2713 (jagr2713):

√x=5+4 Add 4 to 5 to get 9. √x=9 To remove the radical on the left-hand side of the equation, square both sides of the equation. (√x)2=(9)2 Simplify the left-hand side of the equation. x=(9)2

OpenStudy (anonymous):

thanks that really helps too! @jagr2713

OpenStudy (isaiah.feynman):

@jagr2713 Wrong.

jagr2713 (jagr2713):

what i got the answer thou

jagr2713 (jagr2713):

when you simplify you get 81

OpenStudy (isaiah.feynman):

So sqrt of 18 is 9 ?

jagr2713 (jagr2713):

where is 18 in the problem

OpenStudy (anonymous):

9^2 is 81

OpenStudy (anonymous):

answer is 81. @jagr2713 gave a good explanation end of discussion

jagr2713 (jagr2713):

i did 5+4=9^2=81

OpenStudy (isaiah.feynman):

Wrong still. x = 21 The question is \[\sqrt{x-5}=4 \]

jagr2713 (jagr2713):

oh i did √x -5=4

OpenStudy (isaiah.feynman):

@Shay17 might wanna check your work again.

OpenStudy (anonymous):

so its 21?

jagr2713 (jagr2713):

u thought it was a different problem sorry for the mistake

OpenStudy (isaiah.feynman):

Yes.

jagr2713 (jagr2713):

Nice going @Isaiah.Feynman

OpenStudy (anonymous):

thanks! @Isaiah.Feynman